OpenStudy (anonymous):

A population in a certain town is 60% female. You select a 12-person jury from this population at random. What’s the probability that 5 or fewer of the jury members are female?: * a. 0.0573 b. 0.1009 c. 0.1582 d. 0.3348

OpenStudy (anonymous):

Hi, I tried to solve this but I did not get any of the answer choices.

OpenStudy (anonymous):

binomial for this one

OpenStudy (anonymous):

Okay, let me try and solve it with your hint. Please do not give any more help. Thanks!

OpenStudy (anonymous):

k i will shut up

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Okay, binomial distribution says the probability of getting exactly n successes is (n C k)*(p^n)*(p^n-k), where p = probability of success, and n and k are parameters.

OpenStudy (anonymous):

For this problem n =15, and k <= 5.

OpenStudy (anonymous):

So, I believe p(k<=5) = p(k=0) + p(k=1) + p(k=2) + p(k=3) + p(k=4) + p(k=5).

OpenStudy (anonymous):

There might be a faster way to do that...

OpenStudy (anonymous):

Correction, binomial distribution is (n C k)*(p^k)*(p^n-k).

OpenStudy (anonymous):

Okay, here I go.

OpenStudy (anonymous):

p(k=0) = (12 C 0) * (.6)^0*(.4)^12 =0.00001677721

OpenStudy (anonymous):

p(k=1) = (12 C 1) * (.6)^1*(.4)^11 = .000301989888

OpenStudy (anonymous):

p(k=2) = (12 C 2) * (.6)^2*(.4)^10 = .0024914166

OpenStudy (anonymous):

p(k=3) = (12 C 3) * (.6)^3*(.4)^9 = .0124570829

OpenStudy (anonymous):

p(k=4) = (12 C 4) * (.6)^4*(.4)^8 = .0420426547

OpenStudy (anonymous):

p(k=5) = (12 C 5) * (.6)^5*(.4)^7 = .0124570829

OpenStudy (anonymous):

Where (n C k) means "n choose k".

OpenStudy (anonymous):

p(k<=5) = p(k=0) + p(k=1) + p(k=2) + p(k=3) + p(k=4) + p(k=5) = 0.00001677721 + .000301989888 + .0024914166 + .0124570829 + .0420426547 + .0124570829 = 0.06976700419

OpenStudy (anonymous):

Okay, now I need help!

OpenStudy (anonymous):

i can't think of a faster way

OpenStudy (anonymous):

The answer I am getting is not one of the answer choices. Did I make a mistake in my work?

OpenStudy (anonymous):

Correction to the binomial theorem, it should read (n C k)*(p^k)*((1-p)^(n-k))

OpenStudy (anonymous):

\[P(X=k) = \left(\begin{matrix}n \\ k\end{matrix}\right) p^k (1-p)^{n-k}\]

OpenStudy (anonymous):

If they can helps us then they can has a medal.

OpenStudy (anonymous):

sorry it took me a while

OpenStudy (anonymous):

That's alright, let me take a look at what you posted.

OpenStudy (anonymous):

i just added them up

OpenStudy (dan815):

isnt there a better way than to add up all the probability like that

OpenStudy (anonymous):

no women plus one woman plus two women plus 3 women plus 4 women plus 5 women

OpenStudy (anonymous):

if there is, and i am not saying that there isn't, i don't know it

OpenStudy (anonymous):

Was there an error in your input?

OpenStudy (anonymous):

.4^12+12*.6*.4^(11)+66(.6)^2*.4^10+220*.4^3*.6^9+495*.6^4*.4^8+792*.6^5*.4^7

OpenStudy (anonymous):

maybe it is hard to do all of these on the computer

OpenStudy (anonymous):

Should the fourth term be 220*.6^3*.4^9?

OpenStudy (anonymous):

if there is, i don't see it

OpenStudy (dan815):

oh yes cebroski that is right

OpenStudy (anonymous):

Corrected, I think it would look like this: .4^12+12*.6*.4^(11)+66(.6)^2*.4^10+220*.6^3*.4^9+495*.6^4*.4^8+792*.6^5*.4^7

OpenStudy (anonymous):

oh you are right!! i mixed up the .4 and .6!!

OpenStudy (anonymous):

as always, the answer is C

OpenStudy (anonymous):

C for cebroski!

OpenStudy (anonymous):

Thanks y'all for your help. My mistake was that I was not using n C k as a coefficient.

OpenStudy (anonymous):

Now, I understand, the reason we need n C k is because there are n C k ways to choose k women for a 12 person jury.

OpenStudy (anonymous):

I learn good.